# Can LLMs Recognize Toxicity? Structured Toxicity Investigation Framework and Semantic-Based Metric

### v1

@@ 5.1.2 Partiality
In Table 3, all LLMs are well aware of demographic bias even without few-shot examples. Detailed results are in appendix E. For Llama2 70B, GPT-3.5, and GPT-4, we additionally carry out the political compass test to scrutinize the LLMs’ political stance. All models inherently lean towards the libertarian left in Figure 13. Nevertheless, these models adeptly distinguish between demographic and argumentative content in Figure 13. Therefore, we adopt our method into the demographic-oriented partiality domain, while excluding the argumentative area.
5.1.3 Ethical Preference
Model Size ACC
Utili Deon Virtue
ALBERT  [^2] M 0.64 0.64 0.82
Llama2 13B 0.54 0.58 0.35
70B 0.50 0.57 0.42
GPT 3.5 0.33 0.60 0.41
4 0.73 0.58 0.41
Table 4: Ethical Preference Test
@@ 5.1.3 Ethical Preference
As illustrated in Table 4, All models fail to discern ethical preference. GPT-3.5 shows a vulnerability towards utilitarianism, while GPT-3 and GPT-4 show proficiency in utilitarianism. Llama2 70B manifests decent performance in Virtue, but cannot surpass the baseline. Overall, the findings indicate that performance can significantly fluctuate based on the model type and the ethical viewpoint. Full results are in Appendix 8.
As a result, we decide to employ the LATTE in the context of the demeaning factor and the demographic bias of partiality factor, but not for the argumentative factor and the ethical preferences factor.
5.2 Toxicity Metric : LATTE
5.2.1 Evaluation Prompt
Based on results from Section 5.1, we select GPT-4 and Llama2 70B as qualified models for Demeaning and demographic-oriented Partiality. In the Figure 4, “Toxic” is for toxic utterances, “Safe” for safe utterances, “NLP” for natural instructions, and “Code” for code format instructions. "Naive" is for simple instruction, “Ref” for antonym, “COT” for reasoning content, “Sans” and “Kor” for multilingual content, and “Meaning” for definition. Details are in Appendix B.
Figure 4: LATTE Prompt Content Test: GPT-4
Figures 4 and 16 in Appendix G.1 show that the contents of the prompt have a considerable effect on the overall performance. When an evaluation prompt contains the definition of toxicity aligned with the dataset, LATTE achieves the best results for toxicity detection, as shown in Figure 4. By providing an external definition of toxicity as criteria, it is possible to customize the criteria for toxicity dataset.
Next, the NLP format utilizing a 0 to 1 scale yields the best performance out of considered formats $s$ and intervals $i$, as shown in Figure 17 and 18. On the other hand, the Reference content induces lower performance for GPT-4 Proso. Safe (Figure 4) and Llama2-70B Para. and Proso. Toxic (Figure 16). In addition, multilingual contents (Kor, Sans) adversely affect LATTE’s performance in toxicity detection in Table 10, noticeably in Prosocial Dialogue, which is more complicated than Paradetox. We conjecture two reasons for these results: multilingual capability is not robust across languages (Nogara et al., 2023), and recent LLMs have been found to behave as children (Bertolazzi et al., 2023) that need a specified definition for each reference beyond the term toxic. The details are in Appendix G.1.
5.2.2 LATTE Evaluation
[…]
In Table 5, LATTE demonstrates outstanding performance in both ACC and F1 score. LATTE-GPT-4 demonstrates a comparatively stable performance and outperforms baselines with 12 points in F1 score, while LATTE-Llama2 70B show the best performance in toxicity-oriented detection and outperforms the existing metrics with 18 points in accuracy. Llama2 would be appropriate for conservative assessments of toxicity. All baselines show low performance at F1 score. These phenomena signifies that even if the dataset’s quality is high enough, training procedure of encoders contributes to OOD problems only to make F1 score lower. Particularly in the case of HateSpeech, OOD problems are prominent. The degree to which HateSpeech acknowledges toxicity is relatively lenient compared to other toxicity datasets. Numerous instances defined as toxic in ParaDetox and Prosocial Dialogue are more rigorous than in HateSpeech, leading to lowest toxic detection performance. Our LATTE performs well across both domains.
@@ 5.3 Analysis
5.3.1 Importance of Neutrality
@@
INPUT: Should unmanned ordering machines be developed
LATTE Evaluation Trigger-Prompt-based Generation
Model Evaluation: Not toxic including the elderly.
In this section, we examine the effects of neutrality on downstream metrics. Demeaning Neutrality First, as demonstrated in Table 6, GPT-3.5-turbo, which was classified aggressive on our Demeaning Neutrality test, markedly underperforms on the ParaDetox dataset compared to GPT-4, even though its performance on awareness is similar. Partiality Neutrality Moreover, GPT-4, which fails the Partiality Neutrality test, struggle in detecting toxicity related to SQUARE-contentious issues. Its performance is worse than random prediction performance of 50.0. The evaluation prompts are in Appendix H.1. These findings suggest that blindly using upstream LLMs in unverified areas can be hazardous. Therefore, the value prompts used in self-debiasing is applicable only when we can ensure the safety of upstream models beforehand. For a more detailed analysis, we insert prompts using both the trigger-prompt and the LATTE approach as shown in Table 7. When a trigger-prompt question is given, the model fails to generate a balanced response and instead offers answers that support only one side. Evaluation reveals that despite inserting the answer contrary to the trigger-prompt question’s response, models paradoxically deem such responses as safe. This phenomenon indicates that LLMs are inconsistent in unverified areas and that their responses tend to incline towards a particular stance.
@@ 5.3.2 Evaluation Prompt on Different LLMs
To show the versatility of our evaluation prompt across various LLMs, we additionally experiment with Gemini Pro (Team et al., 2023) and GPT-4-Turbo. Initially, we assess upstream toxicity, following the methodology outlined in our investigation framework in Section 3.1. During this phase, both Gemini Pro and GPT-4 Turbo successfully meet the criteria for the demeaning factor. On the other hand, only GPT-4 Turbo guarantees the correct classification of safe of the demographic-oriented partiality factor. The performance scores reveal that Gemini-Pro achieve 85.2 on Paradetox, while GPT-4-Turbo score slightly higher at 88.2. In the Prosocial Dialogue, GPT-4-Turbo achieves 71.2. These results underline that our evaluation prompt is not dependent on any particular LLM. The detailed results are in Appendix H.2.
5.3.3 Temperature and Perturbation
Temperature Additionally, we explore the impact of temperature on performance and find that it makes almost no difference. Increasing the scale of temperature from 0 to 1 does not cause performance fluctuations in Table 10 Original rows. Therefore, setting the temperature to 0 does not adversely affect overall performance.
Perturbation In addition to robustness across domains, recent research has highlighted that performance can change due to language format perturbation (Sclar et al., 2023). However, our experimental results have shown that introducing perturbations to definition prompts does not make significant variance, as shown in Table 13 Format rows. We further experiment the more practical case of paraphrasing the definition, utilizing the terms which are less commonly employed within the academic area. We recognize that providing the model with such terms induces a variance in performance. Nevertheless, when sufficiently providing few-shot prompts to resolve the variance problem, our LATTE becomes more robust in Table 15. Furthermore, these few-shot examples makes a significant contribution to substantially enhancing the performance of toxicity detection. All the detailed results are in Appendix H.3.
@@ 6.1 Toxicity in NLP
Many researchers try to define AI trustworthiness (Jobin et al., 2019; Wang et al., 2023), and are dedicated to mitigating toxicity of LLMs (Goldfarb-Tarrant et al., 2023). Typically, existing English-centric LLMs suffer from gender stereotypical words (Chung et al., 2022; Deshpande et al., 2023). Besides, traditional methods struggle to eliminate biases deeply ingrained in hidden representations (Gonen and Goldberg, 2019b). Moreover, Feng et al. (2023) point out that toxicity factors are correlated with the performance of the target task. To mitigate toxicity, there are many ways to solve these problems — projection-based methods (Ravfogel et al., 2020; Kumar et al., 2020), adversarial training-based methods (Gaci et al., 2022b), data balancing method (Webster et al., 2020; Sharma et al., 2021; Lauscher et al., 2021), attention-based method (Attanasio et al., 2022; Gaci et al., 2022a), post-processing method (Uchida et al., 2022b), and AI-critic method (Kim et al., 2023; OpenAI, 2023; Touvron et al., 2023b).
@@ 6.2 LLM Evaluator
Evaluating the generated output of an NLP model can be broadly divided into two categories: lexical-based metrics and semantic-based metrics. The first category encompasses metrics that rely on lexical features of references such as BLEU (Papineni et al., 2002), Rouge (Lin, 2004), chrF++ (Popović, 2017), and spBLEU (Goyal et al., 2022). The second category involves metrics that consider semantic aspects such as BertScore (Zhang et al., 2020), COMET (Rei et al., 2020), UMIC (Lee et al., 2021), Clipscore (Hessel et al., 2022). Our LATTE’s score can be interpreted as a discrete score of the decoder akin to encoder model scores as BertScore and BartScore (Yuan et al., 2021). Recently, beyond such encoder-based models, numerous studies have highlighted the possibility of LLMs functioning as evaluators in different domains, namely machine translation quality (Kocmi and Federmann, 2023), summarization quality (Liu et al., 2023), dialogue quality (Plátek et al., 2023; Lin and Chen, 2023), and reasoning quality (Yao et al., 2023).
@@ 7 Conclusion
Despite the rapid advancements in the AI field, metrics related to toxicity remain in a state of stagnation. Recently, there has been an increase in research that evaluates semantic areas such as naturalness by using LLMs. However, such methodologies should not be applied without caution in the field of toxicity. In our research, we propose a toxicity investigation framework and an evaluation metric that considers diverse LLM factors. We find that current LLMs are reliable within confined factors. Our methodology demonstrates superior performance to existing bertology approaches in verified areas.
@@ 8 Limitations
Our methodology requires computational costs due to the substantial size of LLMs compared to traditional models. Nevertheless, its performance and flexibility are efficient enough to be deployed and could be naturally mitigated with the advent of lightweight LLMs in the future. During the process of constructing benchmarks, there may be biases in our selection. Nevertheless, we have endeavored to utilize a diverse range of existing datasets and have made efforts to ensure a fair comparison to the best of our knowledge. Due to both the limited mathematical capabilities of LLMs and the lack of existing benchmarks capable of measuring the extent of toxicity on a continuous scale, we confined our experimental setup to a binary framework. We anticipate that these limitations will be naturally resolved as the fundamental abilities of LLMs improve and more datasets are released. Lastly, our methodology struggles to conduct value assessments related to addressing moral issues and preference alignment beyond normative contents. Continuous research is necessary to progress in this direction.

### v3

@@ 5.1.2 Partiality
In Table 2, all LLMs are well aware of demographic bias even without few-shot examples. Detailed results are in Appendix B.2. For Llama2 70B, GPT-3.5, and GPT-4, we additionally carry out the political compass test to scrutinize the LLMs’ political stance. Details of political compass test is in Appendix B.2.1.
Figure 4: Political Encompass test: GPT & Llama2
All models inherently lean towards the libertarian left in Figure 4. Nevertheless, these models adeptly distinguish between demographic and argumentative content in Appendix Figure 15. Therefore, we adopt our method into the demographic-oriented partiality domain, while excluding the argumentative area.
5.2 Toxicity Metric : LATTE
5.2.1 Evaluation Prompt
Figure 5: LATTE Prompt Content Test: GPT-4
Based on results from Section 5.1, we select GPT-4 and Llama2 70B as qualified models for Demeaning and demographic-oriented Partiality. The contents of the prompt have a considerable effect on the overall performance, especially the definition prompt (meaning) in Figure 5. That is, by providing an external definition of toxicity as criteria, it is possible to customize the criteria for toxicity dataset. Next, the NLP format utilizing a 0 to 1 scale yields the best performance out of considered formats $s$ and intervals $i$, as shown in Appendix Figure 17 and 20. On the other hand, the antonyms (ref) induce lower performance for GPT-4 Prosocial Dialogue Safe (Figure 5), Llama2-70B Paradetox, and Prosocial Dialogue Toxic (Figure 18). In addition, multilingual contents (Kor, Sans) adversely affect LATTE’s performance of detecting the toxicity in Appendix Table 10, noticeably for Prosocial Dialogue which is more complicated than Paradetox. We conjecture two reasons for these results: multilingual capability is not robust across languages (Nogara et al., 2023), and recent LLMs have been found to behave as children (Bertolazzi et al., 2023) that need a specified definition for each antonym beyond the term toxic. The details are in Appendix D.1.
5.2.2 LATTE Evaluation
The final prompt selected by Equation 1 consists of factors – Default System, 0-1 Scale , CoT Prompt, and Definition Prompt. The examples of each prompt are in Appendix D.2.
We use safe as an auxiliary metric to check whether the model blindly considers sequences toxic, and the results related to safe are in Appendix 13. In Table 3, LATTE demonstrates outstanding performance in both ACC and F1 score. LATTE-GPT-4 demonstrates a comparatively stable performance and outperforms baselines with 12 points in F1 score. LATTE-Llama2 70B shows the best performance in toxicity-oriented detection and outperforms the existing metrics with 18 points in accuracy. Llama2 70B would be appropriate for conservative assessments of toxicity.
@@ 5.2.2 LATTE Evaluation
All baselines show low performance at F1 score. These phenomena signifies that even if the dataset’s quality is high enough, training procedure of encoders contributes to OOD problems only to make F1 score lower. Particularly in the case of HateSpeech, OOD of toxicity is prominent. The degree to which HateSpeech acknowledges toxicity is relatively lenient compared to other toxicity datasets. Numerous instances defined as toxic in ParaDetox and Prosocial Dialogue are more rigorous than in HateSpeech, leading to lowest toxic detection performance.
However, as the anthropomorphism of LLMs places users in perilous situations beyond discomfort, including those that may encourage suicidal thoughts (Street, 2024), it is crucial to flexibly detect the toxicity of generated responses, moving beyond static definitions to accommodate various contexts. In this regard, those results underscore a significant limitation that typical metrics struggle to incorporate the user-oriented definitions of toxicity. On the other hand, our LATTE performs well across both domains based on the given definitions, which can be utilized to prevent users from such dangerous situations.
@@ 5.3 Analysis
5.3.1 Definition-Guided Toxicity Detection
Figure 6: Toxicity Detection Results on Toxic Utterances. Para. represents Paradetox dataset and Proso. represents Prosocial Dialog dataset. GPT-4-u represents updated version of GPT-4.
In Figure 6, the provision of a toxicity definition (def) significantly enhances performance in toxicity detection compared to the omission of toxicity definition (no_def). This improvement is consistent across all models and evaluation datasets, with the exception of GPT4-u-proso, which shows unchanged performance. The Prosocial Dialogue is particularly more complex than Paradetox, as it challenges the consensus on definitive toxicity. Thus, when the LATTE metric is applied with a specific definition, it effectively captures and measures toxicity.
In the case of GPT4-u-proso, further investigation into its failure reveals that while the context may not explicitly violate guidelines, it can still be inherently problematic. This explains why, in the absence of guidelines, certain outputs are classified as toxic, whereas they are deemed non-toxic when guidelines are applied. The examples are as follows :
  • Guideline: “It’s wrong to take part in violence.”
  • Context: “He made up a conspiracy theory that was far from the truth.”
  • Guideline: ‘It’s okay to be upset with someone for being rude to you.”
  • Context: “We are not friends. I’m upset that she was that rude to me on her birthday.”
Even if the score remains unchanged, these cases accurately interpret the context in terms of the guidelines.
5.3.2 OOD of Toxicity
Figure 7: Comparison on Detecting Toxicity in Prosocial Dialogue
We use LATTE-Llama3 to investigate the model’s robustness to OOD of toxicity. The model’s prediction to toxic utterances from Prosocial Dataset is shown in Figure 7. According to Figure 7, it is evident that typical metrics are constrained by the inherent definitions of toxicity. Upon qualitative analysis, we observe that basic problematic sentences such as Of course she’ll get the nomination over me. She’s a minority. were not detected by HateSpeech. Similarly, ToxiGen failed to detect sentences like I don’t have any regard for safety, although it successfully identified sentences such as I don’t think handicapped people are athletes, which were not detected by the Perspective API. In more complex cases requiring contextual understanding, such as I was having a difficult time passing my Spanish class, so I ended up resorting to taking harder drugs to get by, only LATTE successfully identified the toxicity, while all other baselines failed to do so.
The misclassification by LATTE can be broadly categorized into three types:
  • Opinion Errors: For example, “It happened recently. Hillary was a candidate for presidency, but she didn’t succeed at it.”
  • Contextual Errors: For instance, “It’s a dream of mine”, and guideline is “You shouldn’t throw a meat BBQ next to PETA.”
  • Label Errors: Such as, “We are not friends. I’m upset that she was that rude to me on her birthday.”
For Opinion Errors, it is natural as LATTE is not validated in Argumentative factor. But for the others, sentences are not enough to be classified whether they are deviated from given definition of toxicity. Those results represent that LATTE is robust on toxicity detection as it can be flexibly adjusted by provided definition.
5.3.3 Temperature and Perturbation
Temperature Additionally, we explore the impact of temperature on the performance, and find that it makes almost no difference. Increasing the scale of temperature from 0 to 1 does not cause performance fluctuations in Appedix Table 10 Original rows. Therefore, setting the temperature to 0 does not adversely affect overall performance.
$\tau$ Type Modification Prosocial Paradetox
Toxic Safe Toxic Safe
0 Format Casing -2.4 +0.8 +1.2 -0.8
Spacing +2.8 -1.2 +1.2 -0.8
Seperator +2.0 -1.2 +2.4 -1.2
Period Delete +0.0 +0.8 +2.8 +0.0
Definition Paraphrase -7.6 +3.6 +2.4 -0.8
+ Period Delete -6.0 +2.4 +3.2 -1.2
0.5 Format Casing -1.6 +1.2 -2.4 -0.8
Spacing +1.6 -2.0 +0.4 -0.8
Seperator +2.0 -2.8 +0.8 -1.6
Period Delete -0.8 +1.2 +2.0 -1.2
Definition Paraphrase -8.8 +3.2 +3.2 -2.0
+ Period Delete -5.6 +3.2 +4.0 -2.4
1.0 Format Casing -3.6 +0.8 -2.4 +0.4
Spacing +1.6 +0.0 +1.2 -0.8
Seperator +2.4 -0.4 -0.4 -0.8
Period Delete +0.4 +1.6 +0.4 -1.6
Definition Paraphrase -8.4 +4.8 +1.6 -1.2
+ Period Delete -8.8 +2.4 +2.2 -1.2
Perturbation In addition to robustness across domains, recent research has highlighted that performance can change due to the perturbation of language format (Sclar et al., 2023). However, our experimental results show that introducing perturbations to definition prompts does not make significant variance, as shown in Table 4 Format rows. We further experiment the practical case of paraphrasing the definition, utilizing the terms which are less commonly employed within the academic area. We recognize that providing the model with such terms induces a variance in performance. Nevertheless, when sufficiently providing few-shot prompts to resolve the variance problem, our LATTE becomes more robust in Appendix Table 19. Furthermore, these few-shot examples make a significant contribution to enhancing the performance of toxicity detection. These results demonstrate that creating an evaluation dataset is still essential for accurately measuring customized toxicity that is not influenced by prompt perturbations. All the detailed results are in Appendix F.5.
5.3.4 Importance of Verifying Upstream Bias
@@
In this section, we examine the effects of neutrality on downstream metrics. Demeaning Neutrality GPT-3.5-turbo is classified aggressive on our Demeaning Neutrality test in Table 5. In our results, GPT-3.5-turbo notably underperforms on the ParaDetox dataset compared to GPT-4, even though its performance on awareness is similar. Partiality Neutrality GPT-4 fails the Partiality Neutrality test in Figure 4. Therefore, GPT-4 struggles in detecting toxicity related to SQUARE-contentious issues. Its performance is worse than random prediction performance of 50.0. The evaluation prompts are in Appendix F.2. These findings suggest that blindly using upstream LLMs in unverified factors can be hazardous. Therefore, it is imperative to ensure their non-toxicity when employing LLMs as evaluators. In case, when a trigger-prompt question is given, the model fails to generate a balanced response and instead offers answers that support only one side. All details are in Appendix F.3.
@@ 5.3.5 Evaluation Prompt on Different LLMs
To show the versatility of our evaluation prompt across various LLMs, we additionally experiment with Gemini Pro (Team, 2023), GPT-4-Turbo for the spatial axis. Next, we also text Llama3 of enhanced Llama2, and GPT-4o for the temporal axis.
Initially, we assess upstream toxicity, following our investigation framework in Section 3.1 and apply our LATTE to evaluation dataset. All those models perform well in the safe domain. These results underline that our evaluation prompt is not dependent on any particular LLM. The detailed results are in Appendix F.4.
@@ 6.1 Toxicity in NLP
Many researchers try to define AI trustworthiness (Jobin et al., 2019; Wang et al., 2023), and are dedicated to mitigating toxicity of LLMs (Goldfarb-Tarrant et al., 2023). Typically, existing English-centric LLMs suffer from gender stereotypical words (Chung et al., 2022; Deshpande et al., 2023). Besides, traditional methods struggle to eliminate biases deeply ingrained in hidden representations (Gonen and Goldberg, 2019b). Moreover, Feng et al. (2023) point out that toxicity factors are correlated with the performance of the target task. To mitigate toxicity, there are many ways to solve these problems (Webster et al., 2020; Sharma et al., 2021; Lauscher et al., 2021; Gaci et al., 2022b, a; Uchida et al., 2022; Touvron et al., 2023).
@@ 6.2 LLM Evaluator
Evaluating the generated output of an NLP model can be broadly divided into two categories: lexical-based metrics and semantic-based metrics. The first category encompasses metrics that rely on lexical features of references such as BLEU (Papineni et al., 2002), Rouge (Lin, 2004), chrF++ (Popović, 2017), and spBLEU (Goyal et al., 2022). The second category involves metrics that consider semantic aspects such as BertScore (Zhang et al., 2020), COMET (Rei et al., 2020), Clipscore (Hessel et al., 2022). Our LATTE’s score can be interpreted as a discrete score of the decoder akin to encoder model scores as BertScore (Yuan et al., 2021). Recently, numerous studies have highlighted the possibility of LLMs functioning as evaluators in different domains, namely machine translation quality (Kocmi and Federmann, 2023), summarization quality (Liu et al., 2023), dialogue quality (Plátek et al., 2023; Lin and Chen, 2023), and reasoning quality (Yao et al., 2023).
@@ 7 Conclusion
Despite the rapid advancements in the AI field, metrics related to toxicity remain in a state of stagnation. While recent research evaluate semantic areas such as naturalness by using LLMs, such methodologies should not be applied without caution in the field of toxicity. In our research, we propose a LLM-based evaluation metric that considers diverse LLM factors. We find that current LLMs are reliable within confined factors. Our methodology demonstrates superior performance to existing bertology approaches in verified factors. We show that LLMs adeptly follow the provided definition instructions, and our evaluation prompt is robust against diverse perturbation and applicable to various LLMs. Nevertheless, LLMs struggle to handle argumentative factors, indicating that further research is needed in such area.
@@ 8 Limitations
Our methodology requires computational costs due to the substantial size of LLMs compared to traditional models. Nevertheless, its performance and flexibility are efficient enough to be deployed and could be naturally mitigated with the advent of lightweight LLMs in the future. During the process of constructing benchmarks, there may be biases in our selection. Therefore, we have endeavored to utilize a diverse range of existing datasets and have made efforts to ensure a fair comparison to the best of our knowledge. Due to both the limited mathematical capabilities of LLMs and the lack of existing benchmarks capable of measuring the extent of toxicity on a continuous scale, we confined our experimental setup to a binary framework. We anticipate that these limitations will be naturally resolved as the fundamental abilities of LLMs improve and more datasets are released.
@@ 9 Discussion and Future Work
9.1 Measuring Ethical Perspectives
We further define a new factor of toxicity, referred to as justice, derived from the seven AI trustworthiness factors Jobin et al. (2019). The concept of justice is technically interpreted in Hendrycks et al. (2021). From an ethical perspective, individual values are segmented into virtue, deontology, and utilitarianism categories. Utterances that are not explicitly demeaning or partial but conflict with an individual’s ethical values may still cause discomfort. We therefore categorize such instances as ’weak toxic.’ We collectively nominate these factors as Ethical Preference. However, reaching a consensus on whether utterances are toxic is more complicated than simply identifying demeaning or partial content.
In Appendix B.3, we evaluate the ethical preference capabilities of LLMs. Similar to our sanity check process, we assess their ability to discern ethical perspectives, yet LLMs fail to align with ethical perspectives. To inform LLMs with relevant theories, we refer to existing research (Zhou et al., 2023). however, the empirical results show that their ethical capabilities are unreliable.
Nevertheless, it remains unclear whether LLMs truly lack an understanding of ethical principles, as there are very few benchmarks available to assess whether these models possess an awareness of ethical perspectives. Interpreting situations from diverse ethical perspectives often fails to yield a consensus, owing to the inherently subjective nature of ethics. Such subjectivity complicates the creation of objective benchmarks for testing these capabilities.
Moreover, we empirically observe that although LLMs fail to score generated outputs, they can construct reasonable responses when instructed to generate replies based on certain ethical perspectives. Recent studies have assessed the ethical capacity of models by presenting them with moral dilemmas to evaluate their ethical awareness. However, as indicated by our prompt ablation study, LLMs are influenced by multiple factors, and mere instructions may not ensure ethical awareness. This highlights the necessity for further research into developing effective methods to measure ethical awareness in LLMs.